# Annealed Multiple Choice Learning: Overcoming limitations of Winner-takes-all with annealing

### v1

@@ Appendix C Additional results from UCI Dataset experiments
Experimental setup. The experiments were conducted using the protocol from [28]. Each dataset is divided into 20 train-test folds, except the protein dataset, which had 5 folds, and the Year Prediction MSD dataset which used a single test-set split. We used the same neural network backbone for each baseline: a one-hidden layer MLP with ReLU activation function, containing 50 hidden units except for the Protein and Year datasets, for which 100 hidden units were used. In the WTA models, the final layer of the MLP was duplicated to account for multiple hypotheses $\{f_{k}(x)\}_{k=1}^{n}$ (followed by no activation) and scores $\{\gamma_{k}(x)\}_{k=1}^{n}$ (followed by sigmoid activation). The WTA-based methods (Relaxed-MCL, MCL and aMCL) were trained with $n=5$. All models were trained with the Adam optimizer over 1,000 epochs with a constant learning rate of $0.01$. The data loading pipeline was adapted from the implementation of [26]. The best models for each dataset are highlighted in bold in Table 1 and Table 3, with the second-best models underlined.
Baselines. Table1 of the main paper includes results from three baselines reported from Table 1 of [37], which we use as references for those benchmarks. ‘PBP’ stands for Probabilistic Back Propagation [28], and ‘MC-dropout’ corresponds to Monte Carlo Dropout [20]. Relaxed-MCL was trained with $\varepsilon=0.5$.
@@
Temperature scheduler. For training aMCL in those experiments, a linear temperature scheduler was used, defined as $T(t)=T_{0}(1-\frac{t}{t_{\mathrm{epoch}}})$, where $t$ denotes the current epoch, $t_{\mathrm{epoch}}=1,000$ is the maximum number of epochs, and $T_{0}=0.5$ is the initial temperature.
@@
Table 3: Results on UCI regression benchmark datasets comparing Distortion. Best results are in bold, second bests are underlined.
Distortion ($\downarrow$)
Datasets $\varepsilon$-MCL MCL aMCL
Boston 5.95 $\pm$ 2.94 2.14 $\pm$ 0.48 6.18 $\pm$ 2.99
Concrete 13.69 $\pm$ 3.51 5.13 $\pm$ 1.2 7.97 $\pm$ 3.84
Energy 0.40 $\pm$ 0.13 1.25 $\pm$ 1.22 0.55 $\pm$ 0.51
Kin8nm 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00
Naval 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00
Power 6.59 $\pm$ 0.62 2.26 $\pm$ 0.38 11.93 $\pm$ 2.36
Protein 7.13 $\pm$ 0.14 0.79 $\pm$ 0.02 0.92 $\pm$ 0.05
Wine 0.19 $\pm$ 0.03 0.02 $\pm$ 0.01 0.09 $\pm$ 0.02
Yacht 0.63 $\pm$ 0.28 3.09 $\pm$ 2.35 0.51 $\pm$ 0.37
Year 26.17 $\pm$ NA 4.8 $\pm$ 0.0 4.46 $\pm$ NA
Appendix D Application to speech separation
@@
This Appendix provides a detailed description of the speech source separation experiments. Speech separation consists of isolating the audio signal of each individual speaker from a mixture in which they are simultaneously active. This task is of major interest for speech processing applications such as Automatic Speech Recognition [49, 43], Speaker Diarization [67], or singing voice extraction from music [30]. Most of the recent speech separation systems are deep neural networks trained in a supervised setting, where the ground truth of the separated speaker tracks is known [68].
D.1 Task description
@@
Instead, we propose to use the MCL framework to solve this assignment problem. Naming $\mathcal{L}^{\mathrm{MCL}}_{\mathrm{sep}}$ the resulting loss, and $\mathcal{L}^{\mathrm{aMCL}}_{\mathrm{sep}}$ its annealed variant, we define
@@
$\mathcal{L}^{\mathrm{aMCL}}_{\mathrm{sep}}(\hat{y},y)$ $=\frac{1}{m}\sum_{s=1}^{m}\sum_{k=1}^{n}q_{T(t)}(\hat{y}_{k}\mid y_{s})\ell(\hat{y}_{k},y_{s})\;.$ (38)
@@
D.2 Experimental settings
D.2.1 Separation model architecture
@@
The model configuration used for our experiments is presented in Table 4. It corresponds to the model configuration from the original paper. The variables $l$, $T$ and $P$, which depend on the unknown input length $l$, are not specified.
@@ D.2.2 Training configuration
Training objectives. Three types of systems are compared in the following experiments:
@@
  • aMCL: training with $\mathcal{L}^{\mathrm{aMCL}}_{\mathrm{sep}}$ defined in (38).
@@
Training parameters. The separation models are trained on 5-second audio segments. Since the utterances from the training set might have a longer duration, 5-second segments are randomly cropped inside the training utterances. The utterances shorter than 5s are removed. The batch size is set to 22 to reduce training time compared the standard setup [46, 45, 64], in which a batch size of 4 is used. We verified that our models match the originally reported scores [45] when using this original setting. To analyze robustness to initialization, we train each method with 3 different seeds, and we report the inter-seed average scores and standard deviations. More precisely, the prediction scores are first averaged across the evaluation dataset for each single seed. Then these dataset-wise scores are used to compute the final average and standard deviations. Each model is trained on $t_{\mathrm{epoch}}=200$ epochs, without early stopping. All the training set is processed during each epoch. Unless otherwise stated, the temperature scheduler is chunk linear:
@@
with initial temperature $T_{0}\approx 0.01$ and $t_{\mathrm{max}}=100$. Note that at the end of the training, when $t\in[t_{\mathrm{max}},t_{\mathrm{epoch}}]$, annealing is disabled by setting $T(t)=0$ in order to fine-tune the hypotheses on the target MCL objective.
@@ D.3 Experimental results
  • Appendix D.3.1: comparison of the separation performance of each approach;
  • Appendix D.3.2: impact of the number of hypotheses on the separation performance with MCL and aMCL;
  • Appendix D.3.3: analysis of the phase transitions of aMCL in the context of speech separation.
D.3.1 Separation performance on 2- and 3-speaker mixtures
@@
In the subsequent Sections, we further study these two claims. In Section D.3.2, we demonstrate the advantages of aMCL over PIT by letting the number of predictions $n$ exceed the number of sources $m$. In Section D.3.3, we show that other conclusions of our theoretical analysis concerning phase transitions also seem to hold in the non-Euclidean setting.
Table 5: Comparison of the training method PIT, MCL and aMCL, for the task of source separation of 2 and 3 speakers. The performance is measured using the PIT SI-SDR metric (left) and the MCL SI-SDR metric (right) computed on the WSJ0-mix evaluation subset. Each method is trained using three seeds and we report inter-seed average score and standard deviation.
@@
D.3.2 Impact of the number of hypotheses in MCL and aMCL
In this Section, we study the impact of the number of hypotheses on the source separation performances. We focus on the 3-speaker scenario, and we increase the number of predictions. More precisely, we consider the cases $n\in[3,5,10]$. If $n=3$, the number of hypotheses is the same as the number $m$ of sources to separate. If $n\in[5,10]$ case the number of hypotheses is higher. Since the number of predictions and sources differ, it is not possible to compute the PIT SI-SDR metric, which is ill-defined in this setting. Therefore, we only present results for the MCL SI-SDR metric. In this experiment, we use a single seed across the considered methods to ensure a fair comparison. We report dataset-wise score statistics in Figure 5.
First, we observe that the separation performance increases as the number of predictions $n$ increases, both for MCL and aMCL. This shows that the additional hypotheses are effectively used by both methods, despite the risk of collapse or convergence towards a suboptimal configuration. Recalling that MCL, aMCL, and PIT performances are similar in the case $n=m=3$, this demonstrates a strong advantage of the MCL family compared to PIT. Indeed, we can improve MCL SI-SDR score beyond what is achievable by PIT training, by increasing the number of predictions $n$. Moreover, this improvement comes at minimal computational complexity cost, as discussed in Section D.3.1. This property can be exploited to handle settings where the number of speakers is unknown. This is left to further work.
@@ D.3.2 Impact of the number of hypotheses in MCL and aMCL
Figure 5: Impact of the number of hypotheses. Comparison of MCL (left) and aMCL (right) on the WSJ0-3mix dataset, using the MCL SI-SDR metric. Dataset-wise scores average (blue triangle), median (red line), and quartiles (main box) are reported. The whiskers extend the box by 1.5 times the interquartile gap. The number of hypotheses $n$ is selected among $[3,5,10]$. A higher score indicates better separation performance.
D.3.3 Phase transition in aMCL training

### v3

@@ Appendix D Additional results from UCI Datasets experiments
Experimental setup. The experiments were conducted using the protocol from [29]. Each dataset is divided into 20 train-test folds, except the protein dataset, which had 5 folds, and the Year Prediction MSD dataset which used a single test-set split. We used the same neural network backbone for each baseline: a one-hidden layer MLP with ReLU activation function, containing 50 hidden units except for the Protein and Year datasets, for which 100 hidden units were used. In the WTA models, the final layer of the MLP was duplicated to account for multiple hypotheses $\{f_{k}(x)\}_{k=1}^{n}$ (followed by no activation) and scores $\{\gamma_{k}(x)\}_{k=1}^{n}$ (followed by sigmoid activation). The WTA-based methods (Relaxed-MCL, MCL, and aMCL) were trained with $n=5$. All models were trained with the Adam optimizer over 1,000 epochs with a constant learning rate of $0.01$. The data loading pipeline was adapted from the implementation of [27]. The best models for each dataset are highlighted in bold in Table 2 and Table 1, with the second-best models underlined.
Baselines. Table 2 of the main paper includes results from three baselines reported from Table 1 of [38], which we use as references for those benchmarks. ‘PBP’ stands for Probabilistic Back Propagation [29], and ‘MC-dropout’ corresponds to Monte Carlo Dropout [21]. Relaxed-MCL was trained with $\varepsilon=0.1$ in Tables 1 and 2. Other values of $\varepsilon$ are experimented in Appendix D.2.
@@
Temperature scheduler. For training aMCL in those experiments, an exponential temperature scheduler was used, defined as $T(t)=T_{0}\rho^{t}$, with $\rho=0.95$ and $T_{0}=0.5$, where $t$ denotes the current epoch and $t_{\mathrm{epoch}}=1,000$ is the maximum number of epochs. When the temperature of the exponential scheduler was lower than $45e-$ 4, we set the training back to the vanilla Winner-takes-all mode.
@@
D.1 Impact of the scheduler type in aMCL
The impact of the scheduler type on the performance on the UCI dataset is provided in Table 4. We see that for large dataset sizes, the exponential scheduler outperforms the linear one. Further study on the impact of the speed of the scheduler on the performance is left for further work.
Table 4: Impact of the scheduler type in aMCL. We display the results on the UCI datasets comparing two types of temperature schedules in aMCL. Here, both use an initial temperature $T_{0}=0.5$. aMCL (exp) uses an exponential scheduler of the form $T(t)=T_{0}\rho^{t}$, with $\rho=0.95$, and aMCL (lin) uses a linear scheduler $T(t)=T_{0}(1-\frac{t}{t_{\mathrm{epoch}}})$. The rows are ordered by dataset size $N$.
RMSE ($\downarrow$) Distortion ($\downarrow$)
Datasets aMCL (exp) aMCL (lin) aMCL (exp) aMCL (lin) $N$
Year 9.08 $\pm$ NA 9.20 $\pm$ NA 4.46 $\pm$ NA 4.50 $\pm$ NA 515345
Protein 4.25 $\pm$ 0.02 4.26 $\pm$ 0.04 0.77 $\pm$ 0.03 0.92 $\pm$ 0.06 45730
Naval 8.00e-4 $\pm$ 4.04e-4 1.72e-3 $\pm$ 1.71e-3 5.37e-7 $\pm$ 3.83e-7 3.97e-6 $\pm$ 1.36e-5 11934
Power 4.08 $\pm$ 0.20 4.07 $\pm$ 0.16 2.18 $\pm$ 0.64 13.24 $\pm$ 2.88 9568
Kin8nm 0.08 $\pm$ 0.00 0.08 $\pm$ 0.00 6.81e-4 $\pm$ 8.14e-5 2.94e-3 $\pm$ 1.80e-3 8192
Wine 0.63 $\pm$ 0.04 0.67 $\pm$ 0.05 0.03 $\pm$ 0.01 0.09 $\pm$ 0.02 1599
Concrete 5.47 $\pm$ 0.67 4.99 $\pm$ 0.63 5.71 $\pm$ 1.72 10.18 $\pm$ 4.55 1030
Energy 1.35 $\pm$ 0.97 0.80 $\pm$ 0.33 0.28 $\pm$ 0.09 0.55 $\pm$ 0.51 768
Boston 3.05 $\pm$ 0.91 3.12 $\pm$ 0.68 2.69 $\pm$ 1.39 6.18 $\pm$ 2.99 506
Yacht 1.62 $\pm$ 0.53 0.85 $\pm$ 0.25 1.15 $\pm$ 0.97 0.51 $\pm$ 0.37 308
D.2 Impact of $\varepsilon$ in Relaxed WTA
RMSE compares the barycenter of the predicted distribution with the target positions. For this metric, Relaxed-MCL outperforms other approaches on the UCI datasets when $\varepsilon$ is high (see Table 5 and Figure 6). This outcome is expected, as Relaxed-WTA is biased towards the distribution barycenter under this regime (see Figure 1 of the paper).
However, using RMSE for comparison discards valuable spatial distribution information of the hypotheses. Distortion, defined in (1), addresses this limitation by measuring quantization performance, which our theoretical analysis is based on. When focusing on the Distortion metric, Table 5 shows that lower values of epsilon (e.g., $\varepsilon=0.1$ or annealed epsilon) consistently perform better than for $\varepsilon=0.5$ across nearly all settings. This improvement is attributed to the reduced bias toward the barycenter in this configuration. Additionally, for Distortion, aMCL (trained with an exponential scheduler) generally outperforms Relaxed-WTA on the UCI datasets for the tested Relaxed WTA variants (see Tables 1, 5 and Figure 6). This is especially true on large datasets (Year, Protein).
Overall, aMCL demonstrates a strong balance between Distortion and RMSE compared to the baselines, especially on the largest datasets. Figure 6 further supports this trend, providing additional analysis on Year and Protein, where aMCL, MCL, and Relaxed-MCL are compared across different $\varepsilon$ values in the (RMSE, Distortion) space.
@@ D.2 Impact of $\varepsilon$ in Relaxed WTA
Table 5: Impact of $\varepsilon$ in Relaxed-WTA on UCI regression benchmark datasets comparing RMSE and Distortion. Best results are in bold, second bests are underlined. The annealed version of Relaxed-WTA (R-WTA) uses a linear scheduler starting at $\varepsilon_{0}=0.5$ and decreasing to 0.
RMSE ($\downarrow$) Distortion ($\downarrow$)
R-WTA ($\varepsilon=0.5$) R-WTA ($\varepsilon=0.1$) R-WTA ($\varepsilon$ annealed) R-WTA ($\varepsilon=0.5$) R-WTA ($\varepsilon=0.1$) R-WTA ($\varepsilon$ annealed) $N$
Year 8.91 $\pm$ NA 8.97 $\pm$ NA 9.10 $\pm$ NA 34.65 $\pm$ NA 9.09 $\pm$ NA 4.89 $\pm$ NA 515345
Protein 4.20 $\pm$ 0.03 4.38 $\pm$ 0.02 4.37 $\pm$ 0.05 7.13 $\pm$ 0.16 1.67 $\pm$ 0.16 0.99 $\pm$ 0.07 45730
Naval 1.47e-3 $\pm$ 7.47e-4 1.80e-3 $\pm$ 5.66e-4 2.08e-3 $\pm$ 8.49e-4 1.13e-6 $\pm$ 1.45e-6 4.21e-7 $\pm$ 2.36e-7 7.70e-7 $\pm$ 7.85e-7 11934
Power 3.99 $\pm$ 0.16 4.02 $\pm$ 0.18 4.10 $\pm$ 0.14 7.69 $\pm$ 1.26 2.95 $\pm$ 0.91 4.28 $\pm$ 2.28 9568
Kin8nm 7.68e-2 $\pm$ 1.89e-3 7.82e-2 $\pm$ 1.91e-3 8.64e-2 $\pm$ 5.25e-3 2.82e-3 $\pm$ 2.09e-4 9.32e-4 $\pm$ 7.97e-5 9.06e-4 $\pm$ 3.89e-4 8192
Wine 0.63 $\pm$ 0.04 0.63 $\pm$ 0.04 0.70 $\pm$ 0.06 0.19 $\pm$ 0.03 0.06 $\pm$ 0.02 0.14 $\pm$ 0.05 1599
Concrete 4.91 $\pm$ 0.65 5.28 $\pm$ 0.58 5.14 $\pm$ 0.56 15.08 $\pm$ 5.41 6.91 $\pm$ 2.81 7.40 $\pm$ 3.12 1030
Energy 1.03 $\pm$ 0.21 1.64 $\pm$ 0.36 1.48 $\pm$ 0.59 0.40 $\pm$ 0.13 0.30 $\pm$ 0.12 0.33 $\pm$ 0.10 768
Boston 2.85 $\pm$ 0.57 2.85 $\pm$ 0.72 2.80 $\pm$ 0.69 5.95 $\pm$ 2.94 3.32 $\pm$ 2.84 3.73 $\pm$ 2.38 506
Yacht 1.17 $\pm$ 0.32 2.52 $\pm$ 1.04 1.59 $\pm$ 0.34 0.63 $\pm$ 0.28 1.34 $\pm$ 0.93 0.73 $\pm$ 0.49 308
Figure 6: (RMSE,Distortion) performance on Year and Protein datasets, for Relaxed-WTA with $\varepsilon\in\{0.05,0.1,0.2,0.5\}$ (blue points, with black error bars for std), Relaxed-WTA with annealed $\varepsilon$ (purple) and aMCL (red). Blue lines link the Relaxed-WTA score points. We see that aMCL is below and left of these lines, which indicates a good tradeoff on those two datasets.
Appendix E Application to speech separation
@@
This Appendix provides a detailed description of the speech source separation experiments. Speech separation consists of isolating the audio signals of each individual speaker from a mixture in which they are simultaneously active. This task is of major interest for speech processing applications such as Automatic Speech Recognition [50, 44], Speaker Diarization [71], or singing voice extraction from music [31]. Most of the recent speech separation systems are deep neural networks trained in a supervised setting, where the ground truth of the separated speaker tracks is known [72].
E.1 Task description
@@
Instead, we propose to use the MCL framework to solve this assignment problem. Naming $\mathcal{L}^{\mathrm{MCL}}_{\mathrm{sep}}$ the resulting loss, $\mathcal{L}^{\mathrm{aMCL}}_{\mathrm{sep}}$ its annealed variant, and $\mathcal{L}^{\mathrm{Relaxed-WTA}}_{\mathrm{sep}}$ its relaxed variant, we define
@@
$\mathcal{L}^{\mathrm{aMCL}}_{\mathrm{sep}}(\hat{y},y)$ $=\frac{1}{m}\sum_{s=1}^{m}\sum_{k=1}^{n}q_{T(t)}(\hat{y}_{k}\mid y_{s})\ell(\hat{y}_{k},y_{s})\;,$ (40)
$\mathcal{L}^{\mathrm{Relaxed-WTA}}_{\mathrm{sep}}(\hat{y},y)$ $=\frac{1}{m}\sum_{s=1}^{m}\sum_{k=1}^{n}q_{\varepsilon}(\hat{y}_{k}\mid y_{s})\ell(\hat{y}_{k},y_{s})\;,$ (41)
where $q_{\varepsilon}(\hat{y}_{k}\mid y_{s})=\begin{cases}1-\varepsilon,&\text{if }k=\mathrm{argmin}_{k^{\prime}\in[\![1,n]\!]}\ell(\hat{y}_{k^{\prime}},y_{s})\\ \varepsilon/(n-1),&\text{otherwise}\end{cases}$.
@@
E.2 Experimental settings
E.2.1 Separation model architecture
@@
The model configuration used for our experiments is presented in Table 6. It corresponds to the model configuration from the original paper. The variables $l$, $T$, and $P$, which depend on the unknown input length $l$, are not specified.
@@ E.2.2 Training configuration
Training objectives. Four types of systems are compared in the following experiments:
@@
  • aMCL: training with $\mathcal{L}^{\mathrm{aMCL}}_{\mathrm{sep}}$ defined in (40),
  • Relaxed-WTA: training with $\mathcal{L}^{\mathrm{Relaxed-WTA}}_{\mathrm{sep}}$ defined in (41).
@@
Training parameters. The separation models are trained on 5-second audio segments. Since the utterances from the training set might have a longer duration, 5-second segments are randomly cropped inside the training utterances. The utterances shorter than 5s are removed. The batch size is set to 22 to reduce training time compared to the standard setup [47, 46, 68], in which a batch size of 4 is used. We verified that our models match the originally reported scores [46] when using this original setting. To analyze robustness to initialization, we train each method with 3 different seeds, and we report the inter-seed average scores and standard deviations. More precisely, the prediction scores are first averaged across the evaluation dataset for each single seed. Then these dataset-wise scores are used to compute the final average and standard deviations. Each model is trained on $t_{\mathrm{epoch}}=200$ epochs, without early stopping. All the training set is processed during each epoch. Unless otherwise stated, the temperature scheduler is chunk linear:
@@
with initial temperature $T_{0}\approx 0.1$ and $t_{\mathrm{max}}=100$. Note that at the end of the training, when $t\in[t_{\mathrm{max}},t_{\mathrm{epoch}}]$, annealing is disabled by setting $T(t)=0$ in order to fine-tune the hypotheses on the target MCL objective.
@@ E.3 Experimental results
  • Appendix E.3.1: comparison of the separation performance of each approach;
  • Appendix E.3.2: impact of the number of hypotheses on the separation performance with MCL and aMCL;
  • Appendix E.3.3: analysis of the phase transitions of aMCL in the context of speech separation.
E.3.1 Separation performance on 2- and 3-speaker mixtures
@@
In the subsequent Sections, we further study these two claims. In Section E.3.2, we demonstrate the advantages of aMCL over PIT by letting the number of predictions $n$ exceed the number of sources $m$. In Section E.3.3, we show that other conclusions of our theoretical analysis concerning phase transitions also seem to hold in the non-Euclidean setting.
Table 7: Comparison of the training methods PIT, MCL, and aMCL, for the task of source separation of 2 and 3 speakers. The performance is measured using the PIT SI-SDR metric (left) and the MCL SI-SDR metric (right) computed on the WSJ0-mix evaluation subset. Each method is trained using three seeds and we report inter-seed average score and standard deviation.
@@
E.3.2 Impact of the number of hypotheses in MCL and aMCL
In this Section, we study the impact of the number of hypotheses on the source separation performances. We focus on the 3-speaker scenario, and we increase the number of predictions. More precisely, we consider the cases $n\in\{3,5,10\}$. If $n=3$, the number of hypotheses is the same as the number $m$ of sources to separate. If $n\in\{5,10\}$ case the number of hypotheses is higher. Since the number of predictions and sources differ, it is not possible to compute the PIT SI-SDR metric, which is ill-defined in this setting. Therefore, we only present results for the MCL SI-SDR metric. In this experiment, we use a single seed across the considered methods to ensure a fair comparison. We report dataset-wise score statistics in Figure 7.
First, we observe that the separation performance increases as the number of predictions $n$ increases, both for MCL and aMCL. This shows that the additional hypotheses are effectively used by both methods, despite the risk of collapse or convergence towards a suboptimal configuration. Recalling that MCL, aMCL, and PIT performances are similar in the case $n=m=3$, this demonstrates a strong advantage of the MCL family compared to PIT. Indeed, we can improve MCL SI-SDR score beyond what is achievable by PIT training, by increasing the number of predictions $n$. Moreover, this improvement comes at minimal computational complexity cost, as discussed in Section E.3.1. This property can be exploited to handle settings where the number of speakers is unknown. This is left to further work.
@@ E.3.2 Impact of the number of hypotheses in MCL and aMCL
Figure 7: Impact of the number of hypotheses. Comparison of MCL (left) and aMCL (right) on the WSJ0-3mix dataset, using the MCL SI-SDR metric. Dataset-wise scores average (blue triangle), median (red line), and quartiles (main box) are reported. The whiskers extend the box by 1.5 times the interquartile gap. The number of hypotheses $n$ is selected among $\{3,5,10\}$. A higher score indicates better separation performance.
E.3.3 Phase transition in aMCL training